Leonard, a company that manufactures explosion- proof motors, is considering two alternatives for ex- panding its international export capacity. Option 1 requires equipment purchases of $900,000 now and $560,000 two years from now, with annual M&O costs of $79,000 in years 1 through 10. Option 2 involves subcontracting some of the production at costs of $280,000 per year beginning now through the end of year 10. Neither option will have a sig- nificant salvage value. Use a present worth analysis to determine which option is more attractive at the company’s MARR of 20% per year. (Note: Check out the spreadsheet exercises for new options that Leonard has been offered recently.)

Can you solve it by using excel just
Answer #1

Beginning of year cash for option 2 will be treated as end of year cash for the previous year

using Excel

Year Option 1 Option 2
investment O&M Net Cash flow Net Cash flow
0 -900000 -900000 -280000
1 -79000 -79000 -280000
2 -560000 -79000 -639000 -280000
3 -79000 -79000 -280000
4 -79000 -79000 -280000
5 -79000 -79000 -280000
6 -79000 -79000 -280000
7 -79000 -79000 -280000
8 -79000 -79000 -280000
9 -79000 -79000 -280000
10 -79000 -79000
Net Present Value -1620094 -1408671

As present cost of option 2 is less, it should be selected

Showing formula for excel

Year Option 1 Option 2
investment O&M Net Cash flow Net Cash flow
0 -900000 =I3+J3 -280000
1 -79000 =I4+J4 -280000
2 -560000 -79000 =I5+J5 -280000
3 -79000 =I6+J6 -280000
4 -79000 =I7+J7 -280000
5 -79000 =I8+J8 -280000
6 -79000 =I9+J9 -280000
7 -79000 =I10+J10 -280000
8 -79000 =I11+J11 -280000
9 -79000 =I12+J12 -280000
10 -79000 =I13+J13
Net Present Value =NPV(20%,K4:K13)+K3 =NPV(20%,L4:L12)+L3